# Good News for Polynomial Root-finding

### v35

@@ 6.1 Doubling computational precision towards its a posteriori estimation
Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output error, how should one choose computational precision? A simple practical recipe of [10, 14, 21] was the basis for highly successful performance root-finders of MPSolve: namely, first apply algorithm $\mathbb{A}$ with a lower precision and then recursively double it until an output error decreases below TOL, at which point it outputs precision within a factor of rwo from optimal, thus also supporting near optimal a posteriori bit operation complexity estimate for algorithm $\mathbb{A}$.
In the rest of this section we obtain a priori estimate by proving that precision of $O(b\log(d))$ bits is sufficient for correctness of our subdivision root-finders.
@@ 6.2 Reduction to approximation of $x$NIR$(x)$ within $\frac{3}{8}$
Next we estimate precision of computations in our deterministic and randomized e/i tests that supports approximation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ required in our root-finders; we only need to consider these tests performed for $\theta$-isolated discs $D(c,\rho)$ with $\theta-1>0$ of order $\frac{1}{m}$ where we approximate NIR$(x)$ with the first order divided difference of (1.6).
Map (1.4) reduces such a test for $p(x)$ to the disc $D(0,1)$ for the polynomial $t(x)=p(\frac{x-c}{\rho})$ and decreases the error estimate by a factor of $\rho$, which is at most $2^{b}$ because we only apply our e/i tests to the disc of radius greater than $\epsilon=1/2^{b}$. By examining the impact of map (1.4) on the error and precision bounds we observe that the latter bound can increase by at most a factor of $b$. Thus we arrive at the claimed precision bound $O(b\log(d))$ in case of any disc $D(c,\rho)$ by proving bound $O(\log(d))$ in the case of the unit disc $D(0,1)$.
The output of an e/i test is an integer, and we only need to certify that the output errors of Algs. 4.2 and 4.1 are less than $\frac{1}{2}$.
@@
By slightly abusing notation write ${\rm NIR}(x):=\frac{t^{\prime}(x)}{t(x)}$ rather than ${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}$. According to straightforward error analysis (cf. [20, Appendix A of Ch.3] and the references therein) rounding errors introduced by these additions and division are dominated, and we can ignore them; then it is sufficient to compute $\zeta^{g}{\rm NIR}(\zeta^{g})$ within, say, $\frac{3}{8}$ for every $g$. Indeed, in that case the sum of $q$ error bounds is at most $\frac{3q}{8}$ and decreases below $\frac{1}{2}$ in division by $q$.
6.3 Precision of representation of $x$NIR$(x)$ within $\frac{1}{8}$
@@ 6.4 What is left to estimate?
Instead of $x{\rm NIR}(x)=\frac{xt^{\prime}(x)}{t(x)}$ we actually approximate $\frac{xt^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (1.6). Thus, we shall increase the above error bound $\frac{1}{8}$ by adding to it the upper bounds $\frac{1}{8}$ on $\alpha:=|\frac{xt^{\prime}(x)}{t(x)}-\frac{xt^{\prime}(y)}{t(x)}|$ for $|x|=1$ and on the rounding error $\beta$ of computing $x\frac{t(x)-t(x-\delta)}{\delta t(x)}$. Next we estimate $\alpha$ in terms of $d$, $\theta$, and $\delta$ – by extending the proof of [73, Lemma 3.6], then readily ensure that $\alpha<\frac{1}{8}$ by choosing $\delta$ with a precision $\log_{2}(1/|\delta|)$ of order $\log(d)$, and finally estimate $\beta$. We simplify our estimates by dropping the factor $x$ where $|x|=1$.
@@ 6.5 Precision of approximation with divided difference
Recall the following lemma, which is [73, Fact 3.5].
@@
Under the assumptions of Prop. 6.1 let $w=w(x):=\max_{j=1}\frac{1}{|x-y_{j}|}$. Then
@@
The claimed bound on $|\xi^{\prime}|$ follows from (1.13). It remains to apply the first five lines of the proof of [73, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, s respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
@@
Recall that in our estimates for $\alpha$ and $\delta$ we can apply Prop. 6.1 for $w\leq\frac{\theta}{\theta-1}$ and obtain
@@
Under the assumptions of Prop. 6.1, let $w\leq\frac{\theta}{\theta-1}$. Then
@@ 6.6 The impact of rounding errors
Suppose that a black box oracle computes for us the values $t(y)$ within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 6.2. Then we can ensure that $|\beta|<\frac{1}{8}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $\log(\frac{1}{\nabla(y)})=O(\log(d))$ bits.
@@
By extending Cor. 6.2 obtain $|\alpha^{\prime}|\leq\frac{1}{8}$.
@@
By combining Cors. 6.1 and 6.2 and Prop. 6.3 obtain
@@ 7 Approximation of matrix eigenvalues
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [134, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
@@
This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ (cf. (1.7)) for approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$.
@@
Our root-finders are essentially reduced to evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$. Hence we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
@@
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [19, Eqn. (5)]. [^25] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant. Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion. This computation tends to be more numerically stable than computing determinants and is significantly accelerated in the highly important case where where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
8 Bit operation cost of the classical subdivision root-finding with novel e/i tests based on approximation of a Cauchy integral
@@ 8.1 Our goal and a basic theorem
Next we estimate the bit operation complexity of our root-finders by combining them with fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [40] and Moenck and Borodin [81] (see [18, Sec. 4.5] for historical account), non-trivially extended in [78] to support the following theorem:
@@
For $q=d$ this is [78, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Moroz in [78]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
@@
due to [65, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of the bit operation cost of basic polynomial computations (cf. [130]) with the algorithm of [40, 81]. [^26] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [48, 27, 101, 102] to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (8.1) and (8.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [78, Thm. 2] by applying bound (8.2) to his lower degree polynomials approximating $p(x)$.
@@ 8.2 Reduction of our root-finders to multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ zeros of $p$, we approximate all these zeros within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with deterministic or Las Vegas randomized e/i tests of Sec. 4. Then every subdivision step is reduced to application of e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$ for $\rho_{\lambda}\geq R/2^{b}$ and all $\lambda$.
Such a test amounts to approximation of $p(x)$ at $O(q\log(d))$ equally spaced points $x$ on each of $v$ circles sharing the center $c_{\lambda}$ and having radii in the range $[\rho_{\lambda}/\sigma,\rho_{\lambda}\sigma]$ for $1<\sigma<\sqrt{2}$. Here $v=2m+1$ and $q=m^{2}$ or $v=O(1)$ and $q=m$ in our e/i tests of Sec. 4, while $q=1$ under the random root model.
We only use Thm. 8.1 for $\tau=0$ to cover the case where a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is evaluated at points $x_{j}\in D(0,1)$. To ensure these assumptions we first scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
@@
Due to Cor. 6.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing any $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$. Map the disc $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ into the unit disc $D(0,1)$. This does not change isolation of the boundary circle and may increase the norm $||t(x)||_{1}$ by at most factor of $2^{b+d}$ because $|\bar{c}_{\lambda}|\leq 1$ and $\bar{\rho}_{\lambda}\geq 1/2^{b}$. Now substitute $||t(x)||_{1}\leq 2^{b+d}$ and $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d\log(m))$ (cf. [130, Eqn.(9.4)]) [^28] This bound is readily deduced from the following non-trivial result [130, Thm. 4.2]: Let $t(x)=f\prod_{j=1}^{m}(x-z_{j})\prod_{j=m+1}^{d}(\frac{x}{z_{j}}-1)$, $||t(x)||_{1}=1$, $|z_{j}|<1$ for $j\leq m$, and $|z_{j}|>1$ for $j>m$. Then $|f|\geq 1/2^{d}$. and obtain that $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
@@
For $m=d$ the above bound on $\log(\frac{||t(x)||_{1}}{|t(x)|})$ of [130, Eqn.(9.4)]) is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$ [130, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.

### v36

@@ 10.1 Doubling computational precision towards its a posteriori estimation
Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output errors, how should one choose computational precision? Here is a simple practical recipe from [10, 15, 22] and MPSolve for obtaining a posteriori estimate within a factor of two from optimal: first apply algorithm $\mathbb{A}$ with a lower precision, then recursively double it and stop where an output error decreases below TOL.
In the rest of this section we obtain a priori estimate for the precision of $O(b\log(d))$ bits.
@@ 10.2 Reduction to approximation of $x$NIR$(x)$ within $\frac{3}{8}$
In our e/i tests we approximate NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ with the first order divided difference of (3.2); we only apply the tests to $\theta$-isolated discs $D(c,\rho)$ for $\theta-1>0$ of order $\frac{1}{m}$ and $\rho\geq\epsilon=1/2^{b}$.
Map (1.3) reduces such a test for $p(x)$ to a test in the disc $D(0,1)$ for the polynomial $t(x)=p(\frac{x-c}{\rho})$ replacing $p(x)$ and decreases the error estimate by a factor of $\rho$. Hence map (1.3) increases the precision bound by at most a factor of $b$. Thus we obtain the claimed precision bound in $O(b\log(d))$ in case of any disc $D(c,\rho)$ if we prove bound $O(\log(d))$ in the case of the unit disc $D(0,1)$.
The output of an e/i test is an integer, and we only need to certify that the output error bound of Alg. 9.1 is less than $\frac{1}{2}$. The error can be magnified by a factor of $v$ in $v$ applications of Alg. 9.1 but this factor is immaterial in our notation $\tilde{O}$.
@@
By slightly abusing notation write ${\rm NIR}(x):=\frac{t^{\prime}(x)}{t(x)}$ rather than ${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}$. According to straightforward error analysis (cf. [21, Part A of Ch.3] and the references therein) rounding errors introduced by these additions and division are dominated, and we can ignore them; then it is sufficient to compute $\zeta^{g}{\rm NIR}(\zeta^{g})$ within, say, $\frac{3}{8}$ for every $g$. Indeed, in that case the sum of $q$ error bounds is at most $\frac{3q}{8}$ and decreases below $\frac{1}{2}$ in division by $q$.
10.3 Precision of representation of NIR$(x)$ within $\frac{1}{8}$
@@ 10.4 What is left to estimate?
Instead of $x{\rm NIR}(x)=\frac{xt^{\prime}(x)}{t(x)}$ we actually approximate $\frac{xt^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (3.2). Thus, we shall increase the above error bound $\frac{1}{8}$ by adding to it the upper bounds $\frac{1}{8}$ on $\alpha:=|\frac{xt^{\prime}(x)}{t(x)}-\frac{xt^{\prime}(y)}{t(x)}|$ for $|x|=1$ and on the rounding error $\beta$ of computing $x\frac{t(x)-t(x-\delta)}{\delta t(x)}$. Next we estimate $\alpha$ in terms of $d$, $\theta$, and $\delta$ – by extending the proof of [76, Lemma 3.6], then readily ensure that $\alpha<\frac{1}{8}$ by choosing $\delta$ with a precision $\log_{2}(1/|\delta|)$ of order $\log(d)$, and finally estimate $\beta$. We simplify our estimates by dropping the factor $x$ where $|x|=1$.
@@ 10.5 Precision of approximation with divided difference
Recall the following lemma, which is [76, Fact 3.5].
@@
Under the assumptions of Prop. 10.1 let $w=w(x):=\max_{j=1}\frac{1}{|x-y_{j}|}$. Then
@@
The claimed bound on $|\xi^{\prime}|$ follows from (3.1). It remains to apply the first five lines of the proof of [76, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, s respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
@@
In our estimates for $\alpha$ and $\delta$ we apply Prop. 10.1 for $w\leq\frac{\theta}{\theta-1}$ and obtain
@@
Under the assumptions of Prop. 10.1, let $w\leq\frac{\theta}{\theta-1}$. Then
@@ 10.6 The impact of rounding errors
Suppose that a black box oracle has computed the values $t(y)$ within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 10.2. Then we can ensure that $|\beta|<\frac{1}{8}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $\log(\frac{1}{\nabla(y)})=O(\log(d))$ bits.
@@
By extending Cor. 10.2 obtain $|\alpha^{\prime}|\leq\frac{1}{8}$.
@@
By combining Cors. 10.1 and 10.2 and Prop. 10.3 obtain
@@ 11 Approximation of matrix eigenvalues
Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [138, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
@@
One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using a Las Vegas expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ (cf. (1.6)) for Las Vegas approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$.
@@
Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can apply the following expressions:
@@
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [20, Eqn. (5)]. [^17] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and its determinant. Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion. This computation tends to be more numerically stable than computing determinants and is significantly accelerated in the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
12 Bit operation cost of the classical subdivision root-finding with e/i tests based on approximation of a Cauchy integral
@@ 12.1 Our goal and a basic theorem
Next we estimate the bit operation complexity of our root-finders by combining them with fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [42] and Moenck and Borodin [84] (see [19, Sec. 4.5] for historical account), non-trivially extended in [81] to support the following theorem:
@@
For $q=d$ this is [81, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Moroz in [81]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
@@
due to [68, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of the bit operation cost of basic polynomial computations (cf. [134]) with the algorithm of [42, 84]. [^18] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [50, 28, 105, 106] to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (12.1) and (12.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [81, Thm. 2] by applying bound (12.2) to his lower degree polynomials approximating $p(x)$.
@@ 12.2 Reduction of our root-finders to multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec. 8. Then every subdivision step is reduced to application of e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$ for $\rho_{\lambda}\geq R/2^{b}$ and all $\lambda$.
Such a test, based on our e/i test of Sec. 8, amounts to approximation of $p(x)$ at $O(m\log(d))$ equally spaced points $x$ on each of $v$ circles sharing the center $c_{\lambda}$ and having radii in the range $[\rho_{\lambda}/\sigma,\rho_{\lambda}\sigma]$ for $1<\sigma<\sqrt{2}$.
We only use Thm. 12.1 for $\tau=0$ to cover the case where a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is evaluated at points $x_{j}\in D(0,1)$. To ensure these assumptions we first scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
@@
Due to Cor. 10.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing any $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$. Map the disc $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ into the unit disc $D(0,1)$. This does not change isolation of the boundary circle and may increase the norm $||t(x)||_{1}$ by at most factor of $2^{b+d}$ because $|\bar{c}_{\lambda}|\leq 1$ and $\bar{\rho}_{\lambda}\geq 1/2^{b}$. Now substitute $||t(x)||_{1}\leq 2^{b+d}$ and $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d\log(m))$ (cf. [134, Eqn.(9.4)]) [^20] This bound is readily deduced from the following non-trivial result [134, Thm. 4.2]: Let $t(x)=f\prod_{j=1}^{m}(x-z_{j})\prod_{j=m+1}^{d}(\frac{x}{z_{j}}-1)$, $||t(x)||_{1}=1$, $|z_{j}|<1$ for $j\leq m$, and $|z_{j}|>1$ for $j>m$. Then $|f|\geq 1/2^{d}$. and obtain that $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
@@
For $m=d$ the above bound on $\log(\frac{||t(x)||_{1}}{|t(x)|})$ of [134, Eqn.(9.4)]) is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$ [134, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.